### v4

@@ 2 Random Attention Model with Unobserved References
Let $E\subseteq X$ be a set of reference alternatives. A reference-dependent attention function with respect to $E$ is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
@@
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$ [^8] This is the modified full support assumption, every subset which maintains the attention privilege of the references is paid attention to with positive probability.
The idea here is similar to attention privilege as defined by Kovach and Suleymanov (2023), i.e., if a reference alternative is in the menu, it is always considered. In fact, that is exactly how I define a reference alternative in this setting. A reference alternative can have other properties, which maybe correlated with the preferences, dependent upon some status quo, or some attribute (for instance, the color of a car). However, the presented model is silent on any such dependence. An alternative is a reference if and only if it is attention privileged, irrespective of how it attained such a privilege. In my model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by Kovach and Suleymanov (2023) since the primitives there are menu-reference pairs. Note that the default is only chosen in case none of the alternatives are paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$ and allows us to interpret the default as a reference.
The fundamental analytical task with respect to the presented model is to identify the set of reference alternatives and the underlying preferences. In most canonical papers, it is assumed that the references are observable, which precludes the need to identify the set of references.
@@
The above definition differs from Definition 3 in that it allows for only a reference-dependent attention function. This is the RAM with unobserved references in its most general form.
3 Characterization
In the present section I characterize the RAM-UR by using three testable axioms. These axioms restrict how the stochastic choice function operates across different menus.
Existence of a Dominant Alternative (EDA): For each $A\in\mathcal{X}\setminus\phi$: $p(x,A)=0$ for some $x\in A^{*}$ implies $p(y,\{x,y\})=1$ for some $y\in A\setminus\{x\}$.
A DM with limited attention fails to choose an alternative in two cases. First, if that alternative is never considered. Second, if whenever it is considered, it is dominated by another alternative. EDA says that, the only way an alternative is never chosen is if another alternative always dominates it and is always considered in the binary choice between the two. This means that it is not possible for an alternative to never be chosen just because it is never considered, making each alternative to be non-trivial in the problem. The axiom also says that the set of possible consideration sets is rich enough that the only way an alternative is never chosen is if another dominating alternative is always considered, i.e. it is attention privileged. This way, EDA captures the essence of the modified full support assumption.
@@ 3 Characterization
The three axioms are independent and restrict the stochastic functions only where deterministic choice or non-choice is observed. This is because the model is very flexible on the structure of the attention function, once the full support assumption is satisfied. To further restrict the stochastic function some regularity condition similar to Cattaneo et al. (2020) must be imposed. However, such an imposition is not necessary to get the first set of results described below.
My first result shows that the above three axioms do allow a stochastic choice function to be represented by a RAM-UR model. However, the representation might not be unique.
@@
It is worth mentioning that one can reach multiple representations of a stochastic choice function by slightly changing the constructed attention function. However, in each representation, the set of reference alternatives will be uniquely identified.
@@
The most important identification result is that we are only able to identify the preferences coarsely, i.e. we can identify the relative ranking of an alternative (whether a reference or not) with respect to all reference alternatives and the transitive closure of the partial order constructed through this relative ranking. However, for any two (non-reference) alternatives that have the same relative rankings with respect to all the reference alternatives, we are not able to identify the preference and thus, one can represent the choice function by taking any extension (Szpilrajn (1930)) of the revealed strict partial order [^9] A strict partial order is transitive, irreflexive and asymmetric binary relation defined over $X$.. The coarse identification is the result of the restrictions on both the observability of the references and on the amount of data available.
Even though the axioms restrict choice only where deterministic choice or non-choice is observed, it is not necessary to observe certain choice to identify the set of references. It is also not necessary to observe all menus to make such an identification. The following example illustrates that:
@@
Suppose we have four alternatives $X=\{a,b,c,d\}$ and the following observed choice probabilities:
@@
A natural way to provide a full identification of preferences is to impose some restrictions on the structure of the attention function. First, I will show that the monotonicity condition imposed by Cattaneo et al. (2020) does not offer us any progress in the identification results.
@@
The example below illustrates that even if we restrict ourselves to monotonic attention functions, a unique identification of the preferences is not always possible.
@@ 4.1 Characterization
This axiom is similar to the i-asymmetry axiom that characterizes the IRA (cf. Manzini and Mariotti (2014)). However, it restricts the denominator to be non-zero. Such a restriction was not needed in the IRA characterization, since all probabilities are assumed to be positive in that model. The axiom says that if removing an alternative $b$ changes the choice probability of another, say $a$ (given the positive denominator constraints), then removing $a$ would not change the choice probability of $b$.
@@
Existence of a Dominant Alternative* (EDA*): For all $A\in\mathcal{X}\setminus\phi$: $p(a^{*},A)=0$, implies $p(a^{*},\{x\})=0$ for some $x\in A$.
@@
Regularity (REG): For all $A,B\in\mathcal{X}\setminus\phi$: $A\subset B$ implies $p(a,A)\geq p(a,B)$ for all $a\in A^{*}$.
This is the standard regularity axiom which states that an alternative’s choice probability weakly reduces in a larger menu. This property is in general not satisfied by the RAM-UR. If we consider the IRA model, the i-asymmetry and the i-independence axioms imply the regularity property. However, that is not the case with the other four axioms utilized to characterize the RAM-UR-IRA. Thus, the property is used as an axiom for the characterization. In the absence of non-triviality, this axiom implies that any trivial alternative [^12] An alternative $x\in X$ here is called trivial if $p(x,\{x\})=0$. will not be chosen with a positive probability in any menu.
@@
The proof of Theorem 2 essentially partitions the menus into those with no references and ones with at least one reference. Using the axioms above and the Manzini and Mariotti (2014) result on the set of menus with no reference, it identifies the underlying attention parameter and partial preferences among non reference alternatives (precisely the part of the preferences that were unidentified in the RAM-UR). Since the RAM-UR-IRA is a special case of the RAM-UR, the rest of the identification is easily done similar to the general model. The proof shows that the fitting of the identified preferences through the non-reference menus and the preferences identified through the menus with references does indeed give us a strict total order. Finally, using the first two axioms, we obtain the required functional form.
An advantage of a model like the RAM-UR is that it is in the realm of the stochastic choice function, which is why a nestedness analysis with some classical models can be undertaken. Both the RAM-UR and the RAM-UR-IRA nest rational choice as a special case, i.e. when the set of references is the entire set of alternatives. They both also nest the IRA, i.e. when the structure of the attention function is the IRA structure and the set of references is empty. Finally, the RAM-UR-IRA is also representable by a Random Utility Model (RUM) in a similar spirit to how the IRA can be. This is not true for the RAM-UR since it can violate regularity which is necessary for a RUM representation. A brief discussion on the RUM representation of the RAM-UR-IRA is provided in Appendix 2.
@@ 5 Literature Overview
The idea that limited attention affects behavior and choice is well accepted and widely documented in the literature. Simon’s (1956) model of satisficing choice is one of the earliest models that can be motivated by limited attention. In this model, the heuristic is that the DM stops at the first alternative which is “good enough”. A DM might utilize such a heuristic because of lack of time or attention. Lleras et al. (2017) model “Choice Overload” which leads to worse outcomes in larger menus, the reason being that good options might fail to grab attention in larger sets due to an overload of available options. Masatlioglu et al. (2012) introduce the concept of the attention filter which incorporates the idea that removing an unconsidered alternative does not change the consideration set in a deterministic choice setting where the DM is attention constrained. Dean et al. (2017) is another paper that incorporates limited attention and also the effect of reference alternatives.
@@
Models of stochastic choice ventured into modelling the variability observed in choices due to limited attention. Manzini and Mariotti (2014) was the seminal paper that filled this void. Since then there has been a substantial body of literature combining both limited attention and stochastic choice. A plethora of papers are closely related to Manzini and Mariotti (2014), for instance, Brady and Rehbeck (2016)’s feasibility model is a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility. Another one is Horan (2019) which looks at the IRA without observing the probability of default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and models a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. They discussed the nestedness properties among these special cases. They were unable to conduct the nestedness analysis with respect to the classical stochastic choice models because in their model, the primitives were menu-reference pairs, which placed it outside the realm of stochastic choice models.
@@
The idea of reference-dependent choice has also been the focus of several papers. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^13] Masatlioglu and Ok (2005) provides references for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. They model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They study the existence of a linear order such that the choice from each menu-reference pair is at least as good as the reference point, though not necessarily the maximum [^14] The best alternative according to the preference ordering.. The RAM-UR does have this non-deteriorating property since references are attention privileged, so in any menu, an alternative worse than the reference is never chosen.
@@
We have assumed menu-independence with respect to the references. In Kovach and Suleymanov (2023), the input to the problem is menu-reference pairs so the question of menu independence was moot. However, it is possible to allow for menu-dependent unobservable references as a possible extension to the model in the current paper. This extension obviously has bite, i.e., we cannot explain any stochastic choice function through such a model. For example, any stochastic choice function that allows for alternative $x$ to be chosen with certainty in a menu where $y$ is present and allows alternative $y$ to be chosen with certainty in a menu where $x$ is present cannot be a reference-dependent RAM even with menu-dependent references.
The literature on marketing and consumer research focused on consideration sets and the impact of memory and attention on whether certain brands enter consumers’ consideration sets. Nedungadi (1990) shows that whether an alternative is recalled while forming consideration sets is not necessarily correlated with how highly that alternative is evaluated by the consumer, i.e. attention might be independent of preferences. Advertising often plays a crucial role in increasing an alternative’s possibility of consideration, by increasing its brand recall, or in the terminology of our paper, making it a reference. It is also possible that higher advertising, and thus, more attention, increases the probability of choice of another alternative. This is consistent with both the experiments of Huber et al. (1982) and Nedungadi (1990), and something which can be observed in the model described in this paper. If a reference is added to a menu, the probability of choice of a different alternative can increase. The model described in this paper also seems relevant for more important decision problems like university choice for high school students (Dawes and Brown (2005), Rosen et al. (2008)) and political choice (Wilson (2008)). In these contexts, a fixed set of reference alternatives also makes sense.
Finally, the presence of a default alternative in our model requires some justification. In models of limited attention, the default is used to model the possibility of choice avoidance, delaying choice, which is quite a well-established empirical observation (Dhar (1997)). Horan (2019) argued that it might not be possible to observe the probability of no choice from empirical choice data, and provided a characterization of the IRA by Manzini and Mariotti (2014) without the need for observing such probabilities. However, note that in this model, the default is a de facto reference, and as such, any menu where another reference alternative is present, the default is not relevant. A reference in any menu is the alternative to which any alternative which is considered is always evaluated against. The separation of consideration and evaluation elements of the choice means that the default then becomes the reference for salience. The probability of choosing to abstain is the probability that none of the alternatives will be salient enough to draw attention. If there are other references present, those that are salient enough to always draw attention, the default is redundant.
@@
    We have $xPy$, $yPz$ and $y\in\bar{E}$, so by definition $xPz$ follows.
@@
Given that $P$ is a strict partial order we can extend it to a strict total order [^15] A strict total order is an asymmetric, transitive and complete binary relation over $X$. (Szpilrajn (1930)). Note that such an extension need not be unique. Fix $\succ^{*}$ to be such an extension, that is, for all $x,y\in X$ we have that:
@@
We can have multiple attention functions that satisfy this requirement, I will show one way we can construct a revealed attention function.
@@
Using R-IND and NT, we have $11=\frac{p(a^{*},A\setminus\{a\})}{p(a^{*},A)}=\frac{p(a^{*},\phi)}{p(a^{*},a)}>$ 1
@@
Proof. Suppose not. Then $p(a,\{a,b,c\})=0$. By RIIA, $p(b,\{b,c\})=p(b,\{a,b,c\})>0$. By R-IND, $11=\frac{p(b,\{b,c\})}{p(b,\{a,b,c\})}=\frac{p(b,\{b\})}{p(b,\{a,b\})}>$ 1 which is not possible. So, $p(a,\{a,b,c\})>0$ follows.
@@
Note finally that the RAM-UR might not be a RUM because it fails to satisfy regularity which is a necessary condition for a stochastic choice function to be a RUM, (Falmagne (1978)). This is the case because of the lack of structure on the attention function due to which for an alternative that is undominated by any reference in a smaller and a bigger menu, we can construct attention probabilities of the subsets where it is the best alternative to be more in the bigger menu than in the smaller menu. This will lead to a RAM-UR model that violate regularity. The lack of structure of the general model allows us to construct attention functions that gives us the means to explain different behavioral observations. One such empirically observed behavioral pattern that it allows us to explain is that the status quo (references) is more likely to be chosen in larger sets (Choice Overload) similar to how Kovach and Suleymanov (2023) showed in their model using the Luce Random Attention structure.

### v5

@@ 3 Random Attention Model with Unobserved References: Fixed Set of References
A reference-dependent attention function with respect to a set $E\subseteq X$ of references is a mapping $\mu_{E}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
@@
    $\mu_{E}(D,A)>0$ if and only if $A\cap E\subseteq D\subseteq A$ [^11] This is another version of the modified full support assumption, every subset which maintains the attention privilege of the references is paid attention to with positive probability.
In this model, multiple reference alternatives or no reference alternatives can be present in any menu, both of which are not possible in the model presented by Kovach and Suleymanov (2023) or in GRAM-UR. Note that the default is only chosen in case none of the alternatives is paid attention to, i.e. if $\mu_{E}(\phi,A)>0$. This can only happen if $A\cap E=\phi$ and allows us to interpret the default as the de facto reference.
@@
3.1 Characterization
In the present section I characterize the RAM-UR by using three testable axioms. These axioms restrict the way in which the stochastic choice function operates across different menus.
Existence of a Dominating Alternative (EDA): For each $A\in\mathcal{X}\setminus\phi$: $p(x,A)=0$ for some $x\in A^{*}$ implies $p(y,\{x,y\})=1$ for some $y\in A\setminus\{x\}$.
A DM with limited attention fails to choose an alternative in two cases. First, if that alternative is never considered. Second, if it is considered, and is beaten by another alternative. EDA says that, the only way an alternative is never chosen is if another alternative always dominates it and is always considered in the binary choice between the two. This means that it is not possible for an alternative to never be chosen just because it is never considered, making each alternative to be non-trivial in the problem. The axiom also says that the set of possible consideration sets is rich enough that the only way an alternative is never chosen is if another dominating alternative is always considered, i.e. it is attention privileged. In this way, EDA captures the essence of the modified full support assumption.
@@ 3.1 Characterization
The three axioms are independent, and they impose restrictions on stochastic functions only when deterministic choice or non-choice is observed. This is because the model is very flexible on the structure of the attention function, once the full support assumption is satisfied. To further restrict the stochastic function some regularity condition similar to Cattaneo et al. (2020) must be imposed. However, such an imposition is not necessary to get the first set of results described below.
The result below shows that the above three axioms do allow a stochastic choice function to be represented by a RAM-UR model. However, the representation might not be unique.
@@
It is worth noting that one can obtain multiple representations of a stochastic choice function by slightly modifying the constructed attention function and the extension of the partially identified preferences. However, in each representation, the set of reference alternatives will be uniquely identified, i.e. as compared to the GRAM-UR, the identification of references is better. In the GRAM-UR, the reference identified for each menu was dependent upon the extension of the underlying preference relation. In contrast, here it is independent of it.
@@
Our identification result is that we are only able to identify the preferences coarsely, i.e. we can identify the relative ranking of an alternative (whether a reference or not) with respect to all reference alternatives as well as the transitive closure of the partial order constructed through this relative ranking. However, for any two (non-reference) alternatives that have the same relative rankings with respect to all the reference alternatives, we are not able to identify the preference and thus, one can represent the choice function by taking any extension (Szpilrajn (1930)) of the revealed strict partial order [^12] A strict partial order is a transitive, irreflexive and asymmetric binary relation defined over $X$.. The coarse identification is the result of the restrictions on both the observability of the references and on the amount of data available.
Even though the axioms impose restrictions only when deterministic choice or non-choice is observed, it is not necessary to observe certain choice to identify the set of references. It is also not necessary to observe all menus to make such an identification. The following example illustrates this point:
@@
Consider the set of alternatives $X=\{a,b,c,d\}$ and the following observed choice probabilities:
@@
A natural way to provide a full identification of preferences is to impose some restrictions on the structure of the attention function. First, I will show that the monotonicity condition imposed by Cattaneo et al. (2020) does not lead to better identification results.
@@
The example below illustrates that even if we restrict ourselves to monotonic attention functions, a unique identification of the preferences is not possible.
@@ 4.1 Characterization
This axiom is similar to the i-asymmetry axiom that characterizes the IRA (cf. Manzini and Mariotti (2014)). However, it restricts the denominator to be non-zero. Such a restriction is not needed in the IRA characterization, since all probabilities are assumed to be positive in that model. The axiom says that if removing an alternative $b$ changes the choice probability of another, say $a$ (given the positive denominator constraints), then removing $a$ would not change the choice probability of $b$.
@@
Existence of a Dominating Alternative* (EDA*): For all $A\in\mathcal{X}\setminus\phi$: $p(a^{*},A)=0$, implies $p(a^{*},\{x\})=0$ for some $x\in A$.
@@
Regularity (REG): For all $\in\mathcal{X}\setminus\phi: ABA\subset B$ implies $p(a,A)\geq p(a,B)$ for all $a\in A^{*}$.
This is the standard regularity axiom which states that an alternative’s choice probability weakly reduces in larger menus. This property is in general not satisfied by the RAM-UR. If we consider the IRA model, the i-asymmetry and the i-independence axioms imply the regularity property. However, that is not the case with the other four axioms utilized to characterize the RAM-UR-IRA. Thus, the property is used as an axiom for the characterization. In the absence of non-triviality, REG implies that any trivial alternative [^15] An alternative $x\in X$ here is called trivial if $p(x,\{x\})=0$. will not be chosen with a positive probability in any menu.
@@
The proof of Theorem 3 essentially partitions the menus into menus with no references and such with at least one reference. Using the axioms above and the Manzini and Mariotti (2014) result on the set of menus with no reference, it identifies the underlying attention parameter and partial preferences among non reference alternatives (precisely the part of the preferences that were unidentified in the RAM-UR). Since the RAM-UR-IRA is a special case of the RAM-UR, the rest of the identification is easily done similar to the general model. The proof shows that the fitting of the identified preferences through the non-reference menus and the preferences identified through the menus with references does indeed give us a strict total order. Finally, using the first two axioms, we obtain the required functional form.
An advantage of a model like the GRAM-UR and RAM-UR is that it is in the realm of the stochastic choice function, which is why a nestedness analysis with some classical models can be undertaken. The GRAM-UR, the RAM-UR and the RAM-UR-IRA nest rational choice as a special case. The GRAM-UR, when for each menu, the most preferred alternative is the reference, is the same as rational choice. The RAM-UR and the RAM-UR-IRA are equivalent to rational choice when the set of references is the entire set of alternatives. The RAM-UR and the RAM-UR-IRA also nest the IRA, i.e. when the structure of the attention function is the IRA structure and the set of references is empty. Finally, the RAM-UR-IRA is also representable by a Random Utility Model (RUM) in a similar spirit to how the IRA can be. This is not true for the RAM-UR since it can violate regularity which is necessary for any RUM representation. A brief discussion on the RUM representation of the RAM-UR-IRA is provided in Appendix 2.
The no default restriction on the GRAM-UR prevents it from nesting the RAM-UR. However, with the multiple reference interpretation, restricting how references behave across menus can make these two versions quite close. I discuss the presence of default in attention models and how it might not be a bad assumption in the next section.
@@ 5 Literature Overview
The idea that limited attention affects behavior and choice is well accepted and widely documented in the literature. Simon’s (1956) model of satisficing choice is one of the earliest models that can be interpreted as motivated by limited attention. In this model, the heuristic is that the DM stops at the first alternative which is “good enough”. A DM might utilize such a heuristic because of lack of time or attention. Lleras et al. (2017) modeled choice overload, which leads to worse outcomes in larger menus, the reason being that good options might fail to grab attention in larger sets due to an overload of available options. Masatlioglu et al. (2012) introduce the concept of the attention filter which incorporates the idea that removing an unconsidered alternative does not change the consideration set in a deterministic choice setting where the DM is attention constrained. Dean et al. (2017) is another paper that incorporates limited attention and also the effect of reference alternatives.
@@
Models of stochastic choice ventured into modelling the variability observed in choices due to limited attention. Manzini and Mariotti (2014) is the seminal paper that filled this void. Since then there has been a substantial body of literature combining both limited attention and stochastic choice. A plethora of papers are closely related to Manzini and Mariotti (2014) and include, for instance, Brady and Rehbeck (2016)’s feasibility model as a generalization of the Manzini and Mariotti (2014), though their motivation is not limited attention, but limited feasibility. Horan (2019) looks at the IRA without observing the probability of default. Cattaneo et al. (2020) introduces the general idea of a random attention model. Kovach and Suleymanov (2023) uses the idea of a RAM in the context of reference-dependent choice and models a family of attention functions, each indexed by the reference point, and looked at different special cases like the Constant Random Attention, Luce Random Attention and the Independent Random Attention. They discussed the nestedness properties among these special cases. Their model didn’t lend itself to nestedness analysis with respect to the classical stochastic choice models because in their model, the primitives were menu-reference pairs. This placed it outside the realm of stochastic choice models.
@@
The idea of reference-dependent choice has also been the focus of several papers. Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014) are seminal papers in this area. These authors were one of the first to model status quo bias in choice, which is a widely observed behavior in experiments on individual choice [^16] Masatlioglu and Ok (2005) provides references for such experiments.. Sagi (2006) imposes a no-regret condition on the reference dependent preferences and Tapkı (2007) allows for incomplete reference dependent preferences. Note that all mentioned works model the effect of the status quo on preferences. However, it is possible that the status quo might affect attention as well. Dean et al. (2017) looks at status quo bias and limited attention together, allowing for both reference-dependent preferences and attention. Kovach and Suleymanov (2023), as discussed before, constructs a model where references affect only the attention. Bossert and Sprumont (2009) provides another model where reference points don’t affect preferences. These authors model a deterministic choice function where the choice is consistent with not choosing a worse alternative than the reference. They study the existence of a linear order such that the choice from each menu-reference pair is at least as good as the reference point, though not necessarily the maximum [^17] The best alternative according to the preference ordering.. The GRAM-UR and the RAM-UR have this non-deteriorating property since references are attention privileged, and thus, in any menu, an alternative worse than the reference is never chosen.
@@
The literature on marketing and consumer research focused on consideration sets and the impact of memory and attention on whether certain brands enter consumers’ consideration sets. Nedungadi (1990) shows that whether an alternative is recalled while forming consideration sets is not necessarily correlated with how highly that alternative is evaluated by the consumer, i.e. attention might be independent of preferences. Advertising often plays a crucial role in increasing an alternative’s possibility of consideration, by increasing its brand recall, or in the terminology of our paper, making it a reference. It is also possible that higher advertising, and thus, more attention, increases the probability of choice of another alternative. This is consistent with both the experiments of Huber et al. (1982) and Nedungadi (1990), and is in line with the model described in this paper. If a reference is added to a menu, the probability of choice of a different alternative can increase. The model described in this paper is relevant for more important decision problems like university choice for high school students (Dawes and Brown (2005), Rosen et al. (2008)) and political choice (Wilson (2008)).
Finally, I would like to discuss the incorporation of a default alternative in the proposed framework. In models of limited attention, the default is used to model the possibility of choice avoidance, delaying choice, which is quite a well-established empirical observation (Dhar (1997)). Horan (2019) argued that it might not be possible to observe the probability of no choice from empirical choice data, and provided a characterization of the IRA by Manzini and Mariotti (2014) without the need for observing such probabilities. In a reference-dependent random utility model, the default is a de facto reference, and as such, any menu where another reference alternative is present, the default is not relevant. In the GRAM-UR, the presence of a reference in all menus makes it a standard stochastic choice problem without the possibility of choosing a default. A reference in any menu is the alternative against which any alternative that is considered is always evaluated. The separation of consideration and evaluation elements of the choice means that the default then becomes the reference for salience. The probability of choosing to abstain is the probability that none of the alternatives will be salient enough to draw attention. If there are other references present, those that are salient enough to always draw attention, the default is redundant.
@@
Proof of Theorem 1
Necessity:
ANC: In the GRAM-UR, if an alternative, say $a$ chosen with a positive probability in a menu, while $b$ is not, then that reveals that $a$ is preferred to $b$. A cycle of the form that ANC prevents from existing corresponds with a cycle in this preference. Thus, ANC is necessary.
ND: Since each menu maps to a unique reference alternative in the RAM-UR, the empty set is considered with zero probability. In every other subset, some alternative is the maximal element and, thus, is chosen. So, the sum of choice probabilities of non-default alternatives sum up to one.
Sufficiency:
Define the binary relation $\bar{\succ}$ as follows: $a\bar{\succ}b$ if there exists $A\in\mathcal{X}$ such that $a,b\in A$, $p(a,A)>0$ and $p(b,A)=0$. ANC with $k=2$ ensures that this binary relation is asymmetric, and ANC for arbitrary $k$ ensures that it is acyclic. Define by $\succ$ the transitive closure of $\bar{\succ}$. The above argument shows that $\succ$ is a strict partial order. Note that $\succ$ is the uniquely identified part of the underlying preferences.
Given that $\succ$ is a strict partial order we can extend it to a strict total order [^18] A strict total order is an asymmetric, transitive and complete binary relation over $X$. (Szpilrajn (1930)). Such an extension need not be unique. Fix $\succ^{*}$ to be such an extension, that is, for all $x,y\in X$ we have that:
$x\succ y\implies x\succ^{*}y$
I define the set $D_{x}$ to be the largest subset of $X$ such that $x$ is the $\succ^{*}$-maximal element. Now we will define the revealed attention function for an arbitrary menu. Consider a menu $A\in\mathcal{X}$. Define by $r^{*}_{A}=argmin_{\succ^{*}}\{x\in A:p(x,A)>0\}$, i.e. the $\succ^{*}$-minimal element in $A$ such that it is chosen with a positive probability. We will call this the revealed reference. Note that for a given extension each menu maps to a unique reference.
Now the revealed attention function $\mu^{*}_{r}:\mathcal{X}\times\mathcal{X}$ is specified such that for each $x\in A$:
$p(x,A)=\sum_{D\in\{D:\{r^{*}_{A}\}\subseteq D\subseteq D_{x}\cap A\}}\mu^{*}_{r}(D,A)$
There are multiple ways to construct such an attention function. One way is as follows: For any arbitrary subset such that it does not contain $r^{*}_{A}$, the attention function takes the value zero. For any other subset, some alternative which has a positive choice probability is the maximal element, this is because it contains $r$ and it has a positive choice probability and any alternative which has a positive choice probability is preferred to it according to $\succ^{*}$. This is true because of the choice of $r^{*}_{A}$. We will determine the attention map for such a menu by dividing the choice probabilities of the $\succ^{*}$-maximal alternative among such menus.
Now we can complete the construction similar to that in Kovach and Suleymanov (2023) for an arbitrary alternative $x$ which has a positive choice probability by the following procedure: Let $n(x,A):=|\{D:\{r^{*}_{A}\}\subseteq D\subseteq D_{x}\cap A\}|$. Then we specify $\mu^{*}_{r}(D,A)=\frac{p(x,A)}{n(x,A)}$ for all such $D\in\{D:\{r^{*}_{A}\}\subseteq D\subseteq D_{x}\cap A\}$. Note that this equal division of choice probabilities fulfills the modified full support assumption. Since, the stochastic choice function satisfies ND, the revealed attention function generated will be a generalized reference-dependent attention function according to Definition 4. The binary relation $\succ$ which is the transitive closure of $\bar{\succ}$ gives us the uniquely identified part of the underlying preferences.
@@
    We have $xPy$, yPzyPz and $y\in\bar{E}$, so by definition $xPz$ follows.
@@
Given that $P$ is a strict partial order we can extend it to a strict total order (Szpilrajn (1930)). Note that such an extension need not be unique. Fix $\succ^{*}$ to be such an extension, that is, for all $x,y\in X$ we have that:
@@
We can have multiple attention functions that satisfy this requirement, I will show one way we can construct a revealed attention function similar to how its done in the proof of Theorem 1.
@@
Using R-IND and NT, we have $1=\frac{p(a^{*},A\setminus\{a\})}{p(a^{*},A)}=\frac{p(a^{*},\phi)}{p(a^{*},a)}>1$
@@
Proof. Suppose not. Then $p(a,\{a,b,c\})=0$. By RIIA, $p(b,\{b,c\})=p(b,\{a,b,c\})>0$. By R-IND, $1=\frac{p(b,\{b,c\})}{p(b,\{a,b,c\})}=\frac{p(b,\{b\})}{p(b,\{a,b\})}>1$ which is not possible. So, $p(a,\{a,b,c\})>0$ follows.
@@
Note finally that the GRAM-UR or the RAM-UR might not be RUM because they fail to satisfy regularity which is a necessary condition for a stochastic choice function to be a RUM, (Falmagne (1978)). This is the case because of the lack of structure on the attention function. The lack of structure of the general models allows us to construct attention functions that gives us the means to explain different behavioral observations. One such empirically observed behavioral pattern that it allows us to explain is that the status quo (references) is more likely to be chosen in larger sets (Choice Overload) similar to how Kovach and Suleymanov (2023) showed in their model using the Luce Random Attention structure.